OpenStudy (anonymous):

ball is thrown upward with an initial velocity of 13.0 m/s at an angle of 65.0 above the horizontal Use energy conservation to find the ball's greatest height above the ground

OpenStudy (anonymous):

-mgh=1/2mv′^2-1/2mv^2 v'=13*cos(65°)(m/s) v=13(m/s) h≈7.08(m)